OpenStudy (anonymous):

The given equation involves a power of the variable. Find all real solutions of the equation. x^2=24

myininaya (myininaya):

x=+ or - 24^(1/2)

myininaya (myininaya):

x=+ or -2*6^(1/2)

myininaya (myininaya):

x=2*6^(1/2) x=-2*6^(1/2)

OpenStudy (anonymous):

second one is correct. how do you come to that? it seems as if it's something really simple and i am not connecting it... can you explain?

OpenStudy (anonymous):

i really need help figuring it out, not just the answer.

myininaya (myininaya):

\[x^{2}=9\] wouldn't you agree x=-3 and x=3 since (-3)^2=9 and (3)^2=9

myininaya (myininaya):

?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x^2=24

myininaya (myininaya):

what is (-3)^2?

OpenStudy (anonymous):

-9

myininaya (myininaya):

no (-3)^2=(-3)(-3)=9

myininaya (myininaya):

-3^2=-3*3=9 since the square is only on the 3 and not the whole thing

OpenStudy (anonymous):

i don't understand where the 3 comes from.........

myininaya (myininaya):

x^2=9 just tell what would go in the place of x

myininaya (myininaya):

tell me what you think would go there? ask yourself what is something square that will give me 9?

OpenStudy (anonymous):

i understand. maybe i don't know where you get the 9 then....

myininaya (myininaya):

? i made up the problem

OpenStudy (anonymous):

oh. wtf... lol. i can figure out your problem, why can't i figure out this other one?

myininaya (myininaya):

the 9 just came from somewhere in my head i'm trying to show you something so you understand problems like this

myininaya (myininaya):

lol

OpenStudy (anonymous):

thanks for the example btw :)

myininaya (myininaya):

x^2=9 to solve equations like this you take the square of both sides and you always put + or -

myininaya (myininaya):

the reason is, because as you seen above there are two numbers we can plug into x^2=9 that would be correct just like if we had x^2=25 x=-5 and 5

myininaya (myininaya):

x^2=81 what will your solutions be?

OpenStudy (anonymous):

okay. that makes sense... with your example. i'm guessing because it's very simple is why i can understand it. lol

OpenStudy (anonymous):

x=+or-9

myininaya (myininaya):

you don't like it when the other side isn't a perfect square but its just like these other ones that are you can't get an integer answer though RIGHT! :)

myininaya (myininaya):

x^2=56 what would x be?

OpenStudy (anonymous):

somewhere between 7 and 8.... :/

myininaya (myininaya):

lol x=+sqrt(56) x=-sqrt(56)

OpenStudy (anonymous):

okay. i can write it like that :)

myininaya (myininaya):

you can simplify this x=+ or - sqrt(8*7)=plus or minus sqrt(2*2*2*7)=+ or - 2sqrt(14)

myininaya (myininaya):

like when we did x^2=9 we got x=sqrt(9)=3 and x=-sqrt(9)=-3

OpenStudy (anonymous):

okay. why not 2*sqrt7?

myininaya (myininaya):

sqrt(8*7)=sqrt(2*2*2*7)=sqrt(4*14)=sqrt(4)sqrt(14)=2sqrt(14)

myininaya (myininaya):

sqrt(ab)=sqrt(a)sqrt(b) is a rule you can use in the future k?

OpenStudy (anonymous):

wow. you are very helpful. this is something i am lacking in my knowledge.... :)

myininaya (myininaya):

thanks :)

OpenStudy (anonymous):

give me a sec. i will attempt the original prob...

OpenStudy (anonymous):

wow. i just did it! i can't believe i feel that i actually did it. haha :) x^2=24 x=sqrt+-24 x=sqrt+-(6)sqrt+-(4) x=+-2sqrt(6) look like i'm understanding?

myininaya (myininaya):

very nice but in the steps before don't put the +- inside the sqrt x=+-sqrt(24) x=+-sqrt(6)sqrt(4) x=+-2sqrt(6) but very nice otherwise :) gj

myininaya (myininaya):

can you do x^2=54

OpenStudy (anonymous):

i see. sorry. having internet issues here. :/

OpenStudy (anonymous):

x=+or- 3sqrt(6)?

OpenStudy (anonymous):

i believe this makes sense. :D i would like to know if the last one i did is correct :)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x^2%3D54 MHUAAHAHA!!! You are the best!